# Nested Sampling for ARIMA Model Selection in Astronomical Time-Series Analysis

### v1

@@ 4.3 Kepler and TESS lightcurves
Figure 15: Autocorrelation and Partial autocorrelation plots for the lightcurve of KIC 12008916 (top) and the residuals (bottom) obtained after subtracting the best fit ARIMA(0,0,1)(0,0,1) model from the lightcurve data. The model fit has been able to capture most of the autocorrelation in the lightcurve data.
The ARIMA(0,0,1)(0,0,1) model has therefore successfully captured the stochastic variability in the original lightcurve.
Ross 176
The detection of transiting exoplanets is sometimes impeded due to the presence of autocorrelated noise in the lightcurves (Pont et al., 2006). This autocorrelated variability may arise from activity of the host star, instrumental artefacts, or other unexplained effects. Autoregressive modelling in such cases is effective in removing the correlated noise and enhancing the signal-to-noise ratio of transit detections (Melton et al., 2024). We demonstrate the results of applying the ARIMA-Nested Sampling framework to the photometric lightcurve data of Ross 176 - a late K-type main-sequence star. The star was recently confirmed to host a transiting super-earth exoplanet - Ross 176b (Geraldía-González, S. et al., 2025), with a period of approximately $5$ days. For this analysis, we used the PDCSAP flux of the TESS photometric lightcurve, corresponding to the sector 83 and year 2024 (Figure 16). The lightcurve was obtained using the MIT quick-look pipeline (Huang et al., 2020a, b).
Figure 16: Normalized TESS photometric lightcurve of Ross 176 from Sector-83, processed using the MIT Quick Look Pipeline (QLP). Dashed red lines indicate gaps in the data where we perform the partition into the corresponding training datasets and label them numerically.
The data was partitioned into five training datasets corresponding to the presence of five significant gaps in the total time series data. The fourth and fifth training datasets exhibit an obvious trend. This was further confirmed by the ADF and KPSS tests performed on these datasets, which indicated non-stationarity. They are therefore subjected to first order differencing ($d=1$) before further analysis. All datasets showed some level of autocorrelation, with datasets (1)(1),$(4)$ and $(5)$ depicting significant autocorrelation, even after differencing (see Figure 17).
We performed an ARIMA grid search on the five training datasets, limiting the maximum ARMA orders to $p=q=4$. For the last two training datasets, the sampler failed to converge for higher orders. We present in Table LABEL:model_log_post_table , the model log posterior probabilities for these datasets calculated up to the maximum possible ARMA orders before the sampler failed to converge.
Table 5: Log posterior probabilities and their associated uncertainties for ARIMA(p,1,q)(p,1,q) model fit to training datasets (4) and (5) in Figure 16.
$\mathbf{(p,1,q)}$ Dataset (4) Dataset (5)
$\log{P_{i}}$ $\sigma_{\log{P_{i}}}$ $\log{P_{i}}$ $\sigma_{\log{P_{i}}}$
(0,1,1) $7.7228\times 10^{-5}$ 0.7672 $1.7548\times 10^{-6}$ 0.7179
(0,1,2) $9.4965\times 10^{0}$ 0.8216 $1.4139\times 10^{1}$ 0.8151
(0,1,3) $1.3768\times 10^{1}$ 0.8518 $1.5256\times 10^{1}$ 0.9782
(0,1,4) $1.7617\times 10^{1}$ 0.8893 $1.7444\times 10^{1}$ 0.9257
(1,1,0) $4.7834\times 10^{2}$ 0.6988 $5.3737\times 10^{2}$ 0.7782
(1,1,1) $1.7541\times 10^{1}$ 0.9768 $2.2808\times 10^{1}$ 0.9227
(1,1,2) $2.2764\times 10^{1}$ 0.9711 $3.7526\times 10^{1}$ 0.9353
(1,1,3) $1.6884\times 10^{1}$ 1.0118 $2.1302\times 10^{1}$ 1.0024
(1,1,4) $3.5210\times 10^{1}$ 0.9806 $2.1779\times 10^{1}$ 0.9704
(2,1,0) $3.3926\times 10^{2}$ 0.8071 $3.4895\times 10^{2}$ 0.8189
(2,1,1) $1.5017\times 10^{1}$ 0.8937 $1.4081\times 10^{1}$ 0.9881
(2,1,2) $1.4214\times 10^{1}$ 0.9279 $9.4836\times 10^{1}$ 1.2746
(2,1,3) $4.5073\times 10^{1}$ 1.2798 $3.4356\times 10^{1}$ 0.9419
(2,1,4) $3.4459\times 10^{1}$ 0.7911
(3,1,0) $2.8067\times 10^{2}$ 0.8213
(3,1,1) $3.1983\times 10^{1}$ 1.0630
(3,1,2) $3.9954\times 10^{2}$ 1.0939
(3,1,3) $4.7549\times 10^{1}$ 1.0840
From the results of ARIMA grid search and Table LABEL:model_log_post_table, the best fit ARIMA models for the training datasets $(1)$, $(2)$, $(3)$, $(4)$ and $(5)$ were identified to be ARIMA(1,0,2)(1,0,2), ARIMA(1,0,0)(1,0,0), ARIMA(2,0,2)(2,0,2), ARIMA(0,0,1)(0,0,1) and ARIMA(0,0,1)(0,0,1), respectively. These models were fit to the training dataset using a high resolution nested sampling run. The posterior means were used to generate forecasts for the training datasets and obtain the time ordered residuals. The right panel (b) of Figure 17 show the autocorrelation and partial autocorrelation plots of the residuals for datasets $(1)$, $(4)$ and $(5)$. The lack of significant autocorrelation in these residuals, as seen from the plots, show that the best fit ARIMA models have successfully captured the variability of the datasets.
Figure 17: Autocorrelation function (ACF) and Partial Autocorrelation function (PACF) plots for the training datasets $(1)$, $(4)$, and $(5)$ (from top to bottom) of Ross 176 (see Figure 16) in the left column (a), and for the residuals, obtained from the best fit ARIMA model, in the right column (b). The lightcurves exhibit strong autocorrelation, which is effectively removed using ARIMA modelling, as demonstrated by the near-white ACF/PACF structure in the residuals.
@@ 5 Conclusions
In this paper, we presented a novel method for ARIMA model selection using a fully vectorized nested sampling algorithm. The framework was first validated on an artificial AR(2)(2) and ARIMA(1,1,1)(1,1,1) process, where it successfully recovered the true underlying model order, and subsequently, posteriors centred on the true values of the model parameters. Its application to sunspots number data indicated preference for higher order ARIMA models. This can be explained by the long-memory behaviour and variability across multiple timescales of the solar magnetic activity. These additional autocorrelation structures cannot be captured by low-order models, and the Bayesian evidence therefore justifies the increased complexity. This highlights the ability of our Nested Sampling framework to identify genuinely multi-scale stochastic processes without overfitting. The data was fit to an ARIMA(9,0,1)(9,0,1) model which yielded accurate multi-step future forecasts using the predictive posterior functions. Finally, the framework was applied for analysing photometric lightcurves of KIC 12008916 and Ross 176 using Kepler and TESS data, respectively. The ARIMA models, selected from the nested sampling runs, were able to completely capture the autocorrelation present in both datasets.
@@
A significant computational bottleneck in this work was sampling valid initial points for higher order ARMA coefficients $\phi_{a}$ and $\theta_{m}$ using the process discussed in Subsection 3.1. If very strong autocorrelation in the time-series is not expected a priori, then the “rejection sampling" process can be sped up by regularizing and constricting the priors using $\sigma^{\prime\prime}<1$ (Table 1). A more powerful approach is offered by reparametrization of the ARMA(p,q)(p,q) coefficients to a set of $p+q$ partial autocorrelation coefficients (Barndorff-Nielsen & Schou, 1973). These coefficients are naturally constrained between 1-1 and $1$, and therefore a simple uniform prior, bounded between these two values, can be implemented.
@@
We note that future improvements to the multi-step forecasts could be made by adopting a rolling window, which is a more accurate (although computationally expensive in our case) way of using ARIMA forecasting (Ndungi & Stanislavovich, 2025). Additionally, a potential forecasting refinement within a Bayesian framework involves a short time-step rolling forecast and using the obtained posterior distributions as priors for the next forecast iteration. Our analysis in this work, though limited to classic ARIMA models, can also be effectively applied to its hybrid extensions, such as - Seasonal ARIMA (SARIMA), Seasonal ARIMA with exogenous variables (SARIMAX), Continous ARIMA (CARIMA) and Autoregressive Fractionally Integrated Moving Average (ARFIMA) models (Feigelson et al., 2018). These extensions can be more effective in modelling both periodic and stochastic variability exhibited in many astronomical time series, such as those of transiting exoplanets, eclipsing binaries or Active Galactic Nuclei (AGN).
@@
The investigation of potential physical interpretations behind the use of autoregressive modelling in time domain astronomy forms another compelling future direction of work. This question could be explored by applying our model selection procedure to different time-varying astronomical systems. If systems of a particular class tend to favour similar model structures then the ARIMA model may be representative of an underlying physical process rather than being mere statistical descriptions of the data.

### v2

@@ 4.3 Kepler and TESS light curves
Figure 17: Autocorrelation (ACF) and Partial autocorrelation (PACF) plots for the light curve of KIC 12008916 (top) and the residuals (bottom) obtained after subtracting the best fit ARIMA(0,0,4)(0,0,4) model from the light curve data. The residual ACF/PACF plots show that the model fit has reduced most of the correlation present in the light curve data.
The results of the Ljung-Box test, up to lag $10$, on the time-ordered residuals (see Table 6) also show no correlation at the 5% significance level. The ARIMA(0,0,4)(0,0,4) model has therefore successfully modelled the stochastic variability in the original light curve of KIC 12008916.
Table 6: Ljung–Box (LB) test p-values for residual autocorrelation of KIC 12008916 data.
Lag p-value
1 0.720115
2 0.906690
3 0.765711
4 0.871563
5 0.936565
6 0.963805
7 0.539844
8 0.602335
9 0.667972
10 0.540848
Kepler 17
The detection of transiting exoplanets is sometimes impeded due to the presence of autocorrelated noise in the light curves (Pont et al., 2006). This autocorrelated variability can arise from activity of the host star or also from systematics. Autoregressive modelling of the correlated noise is effective in such cases for enhancing the signal-to-noise ratio of transit detections (Melton et al., 2024).
Here, we demonstrate a proof-of-concept by applying the ARIMA-Nested Sampling framework to the photometric light curve data of Kepler 17 – a very active, main-sequence star of the spectral class G. The star was confirmed to host a transiting hot Jupiter – Kepler 17b (Désert et al., 2011), with a period of $1.5$ days (frequency $\approx 0.66/\text{day}$). For this analysis, we used the normalized PDCSAP flux of the Kepler light curve, corresponding to the Kepler Quarter 16 and year 2013 (Figure 18). The original light curve had a sampling gap of around $28$ minutes in the data while the cadence was around $0.86$ minutes. Figure 18 was obtained by binning the data into time-bins of $0.0211$ days ($30.4$ minutes), to render the sampling rate uniform. By doing so, we focus on the stellar variability exhibited on longer scales but lose information about the short-timescale stochastic variability.
Figure 18: The normalized photometric light curve of Kepler-17, from Kepler Quarter 16, showing periodic transit dips caused by the planet Kepler-17b. The transit dips are superimposed on a quasi-periodic flux resulting primarily from spot modulation – a combined effect of starspots and stellar rotation.
This light curve exhibits a significant transit-depth from the exoplanet and spot modulation due to the stellar activity which also reflects the rotation rate of the star (Lanza, A. F. et al., 2019). The data was tested for stationarity using the ADF and KPSS tests, and was found to be non-stationary according to the ADF test (see Table 7). The time series was rendered stationary after first-order differencing $d=1$.
Table 7: Results of Augmented Dickey-Fuller (ADF) and KPSS Stationarity Tests for Kepler 17 light curve.
Statistic ADF Test KPSS Test
Test Statistic -2.438782 0.431541
p-value 0.131088 0.063560
Lags Used 14 24
Number of Observations Used 1493
Critical Value (1%) -3.434738 0.739000
Critical Value (5%) -2.863478 0.463000
Critical Value (10%) -2.567802 0.347000
The grid search was performed on this differenced data up to a maximum AR and MA order of $7$. The results of this grid search, shown in Figure 19, reveal ARIMA(1,1,5)(1,1,5) as the model with the highest log-posterior probability ($\log{P_{\text{max}}}=-0.0054\pm 0.88$).
Figure 19: Heatmap of the ARIMA model’s log posterior probabilities obtained from the nested sampling runs on the Kepler 17 data (Figure 18). The differencing order is fixed to $d=1$. The grid search shows ARIMA(1,1,5)(1,1,5) as the model with highest log posterior probability.
A high-resolution nested sampling run ($n_{\text{live}}=500$) was performed to fit the ARIMA(1,1,5)(1,1,5) model to the data, and as before, time-ordered residuals using posterior means were obtained.
Figure 20: Results of the ARIMA(1,1,5)(1,1,5) model fit to the normalized light curve of Kepler 17 (Figure 18). Top panel shows the light curve data (in black) and the model fit curve (in red) obtained from the mean of the posterior samples. Bottom panel shows the corresponding residual time series. The residuals depict clear periodic spikes corresponding to the transit dips of Kepler 17-b.
The residual time series still depicts a periodic structure resulting from the planet transits. This shows that even though we have not masked the transits during the fit, the ARIMA model comparison and fit is not completely biased by the transits. For brevity, we will not show the histogram of these residuals, but inspect the ACF/PACF plots as shown in Figure 21.
Figure 21: Autocorrelation and Partial autocorrelation plots for the light curve of Kepler 17 (top) and the residuals (bottom) obtained after subtracting the best fit ARIMA(1,1,5)(1,1,5) model from the light curve data. The residual ACF/PACF plots show that the model fit has reduced most of the correlation present in the light curve data.
These plots show that the model has removed most of the correlation present in the original data. Table 8 shows results of the Ljung–Box test on the residuals. These results show statistically significant autocorrelation at lags 2, 3, and 5, indicating that the ARIMA(1,1,5) model has not fully whitened the residuals in the strict sense required for a complete ARIMA fit. This residual correlation is, however, physically expected. The periodic transit signature of Kepler-17b at a period of approximately 1.5 days introduces a deterministic, non-stochastic signal component that ARIMA, as a purely stochastic model, is not designed to absorb. A more complete, future treatment would involve masking the transit epochs prior to ARIMA fitting, or employing a hybrid model that explicitly accounts for both the stochastic stellar component and the periodic planetary signal.
Table 8: Ljung–Box (LB) test p-values for residual autocorrelation of Kepler 17 light curve.
Lag p-value
1 0.116686
2 0.021184
3 0.044350
4 0.067597
5 0.029983
6 0.051579
7 0.079548
8 0.091245
9 0.103069
10 0.121856
To further determine whether this fit has removed the periodic signal due to stellar activity, we inspect the periodograms of the light curve and the residuals in Figure 22. Figure 22 shows the Lomb-Scargle periodograms of the light curve data and the residual from the ARIMA model fit, evaluated on a frequency grid of $0.001/\text{day}$ to $1/\text{day}$.
Figure 22: Lomb-Scargle periodograms of the light curve data (black) and ARIMA(1,1,5)(1,1,5) fit residuals (blue) from Figure 20. The periodogram of the original light curve shows multiple peaks associated with stellar variability, with the strongest peak corresponding to the stellar rotation period of $\sim 11$ days. The peak near $0.66/\mathrm{day}$ corresponds to the transit period of Kepler-17b ($\sim 1.5$ days). In the residual periodogram, the stellar variability is significantly suppressed while the planetary transit signal remains relatively prominent.
The light curve periodogram has many peaks at different frequencies. The period at maximum power corresponds to the known stellar rotation period of approximately $11$ days. The peak at frequency corresponding to $1.5$ days, marked by the red dashed line, is from the periodic transit signals of Kepler 17b. We estimate the relative prominence of the transit signal by computing the ratio between the periodogram power at the transit frequency and the power at the stellar rotation frequency. This ratio from the light curve periodogram was calculated to be around $0.21$. For the residual’s periodogram, the period at maximum power corresponded to the transit period of $1.5$ days. A similar calculation of the relative strength of the transit signal with respect to the power at the stellar rotation peak yielded a value of $1.27$. This ratio increase from $0.21$ to $1.27$ indicates that the ARIMA(1,1,5)(1,1,5) fit has modelled the stellar rotation signal while relatively preserving the transit periodicity.
3C 273 and S4 0954+65
Quasars represent the brightest and most extreme objects in the universe and fall under the broader classification of Active Galactic Nuclei (AGN). The photometric variability of AGNs in multiple wavelengths is an active area of research. The study of their variability offers insights into the physical processes governing accretion and the activity of the AGN’s central engine. ARIMA can be used to effectively model the stochastic variability depicted by the light curves of such quasars (Hanif & Protopapas, 2015; Belete et al., 2019).
Here, we present the results of applying our ARIMA-Nested Sampling framework to two quasars – 3C 273 and S4 0954+65. The former was identified in the Third Cambridge Catalogue of Radio Sources (Archer et al., 1959). It is the brightest quasar in apparent magnitude and the first to be spectroscopically identified. S4 0954+65 is a blazar of the BL Lac subtype, discovered in the S4 radio survey (Cohen et al., 1977). Both quasars were extensively observed by the TESS mission over multi-day observing sectors. In this analysis, we use data from the Sector 46 (year 2021) for 3C 273 and Sector 20 (year 2019) for S4 0954+65. The data was obtained and processed using the MIT Quick Look Pipeline (QLP) (Huang et al., 2020a, b) for 3C 273 and using the TESS Science Processing Operations Center (TESS-SPOC) pipeline (Jenkins et al., 2016) for S4 0954+65.
Figure 23 shows the normalized light curves of 3C 273 and S4 0954+65. The latter is produced by binning the original flux data into time bins of $20.16$ minutes while the cadence of the original light curve data was around $1.87$ minutes.
Figure 23: TESS normalized light curves of the quasar 3C 273 (left) and the blazar S4 0954+65 (right) plotted on a shared y-axis. S4 0954+65 light curve is obtained by binning the original light curve in time with a bin size of $20.16$ minutes. From the plots, S4 0954+65 exhibits significantly more optical variability than 3C 273.
There is a single significant sampling gap in both light curves. Therefore, for the grid search and rest of the analysis, we use the data before the sampling gaps as our training dataset. The light curve of 3C 273 shows little to no variability compared to S4 0954+65, except the two jitters due to pointing instabilities. These bad values are also clipped from our training data. The results of the ADF and KPSS stationarity tests (see Table 9 and Table 10) indicated both time series to be non-stationary. This non-stationarity is also apparent from their light curves, where the 3C 273 light curve exhibits a linear trend and S4 0954+65 light curve shows heteroscedasticity.
Table 9: Results of Augmented Dickey-Fuller (ADF) and KPSS Stationarity Tests for 3C 273 light curve.
Statistic ADF Test KPSS Test
Test Statistic -1.623081 5.853438
p-value 0.471114 0.010000
Lags Used 23 26
Number of Observations Used 1586
Critical Value (1%) -3.434480 0.739000
Critical Value (5%) -2.863364 0.463000
Critical Value (10%) -2.567741 0.347000
Table 10: Results of Augmented Dickey-Fuller (ADF) and KPSS Stationarity Tests for S4 0954+65 light curve.
Statistic ADF Test KPSS Test
Test Statistic -3.971632 0.514357
p-value 0.001567 0.038433
Lags Used 4 18
Number of Observations Used 845
Critical Value (1%) -3.438112 0.739000
Critical Value (5%) -2.864966 0.463000
Critical Value (10%) -2.568595 0.347000
We therefore fix the differencing order to $d=1$ for the grid searches. The results of the grid search for both light curves are shown in Figure 24 and Figure 25.
Figure 24: Heatmap of the ARIMA model’s log posterior probabilities obtained from the nested sampling runs on the 3C 273 light curve data. The differencing order is fixed to $d=1$. The grid search shows ARIMA(1,1,1)(1,1,1) as the best-fit model with highest log posterior probability.
Figure 25: Heatmap of the ARIMA model’s log posterior probabilities obtained from the nested sampling runs on the S4 0954+65 light curve data. The differencing order is fixed to $d=1$. The grid search shows ARIMA(1,1,4)(1,1,4) as the best-fit model with highest log posterior probability.
From the grid searches, ARIMA(1,1,1)(1,1,1) is picked as the model with highest log posterior probability ($\log{P_{\text{max}}}=-0.0243\pm 0.79$) for 3C 273 data and ARIMA(1,1,4)(1,1,4) ($\log{P_{\text{max}}}=-1.1964\pm 0.81$) for S4 0954+65. We therefore fit these models to the data with a high-resolution nested sampling run, and show the results in Figures 26 and 27.
Figure 26: Results of the ARIMA(1,1,1)(1,1,1) model fit to the normalized light curve of 3C 273 (Figure 23).
Figure 27: Results of the ARIMA(1,1,4)(1,1,4) model fit to the normalized light curve of S4 0954+65 (Figure 23).
The residual time series for both 3C 273 and S4 0954+65 show no temporal structure. Furthermore, the histograms of these residuals, as shown in Figure 28, appear to be normally distributed around 0.
Figure 28: Histograms of the residuals of 3C 273 (left) and S4 0954+65 (right) obtained after subtracting the best fit ARIMA models. The residuals appear to be normally distributed around zero.
The ACF and PACF plots of the light curves and residuals (see Figures 29 and 30) show that the respective ARIMA model fits have successfully removed the correlation in the original light curves. The results of the Ljung-Box test on both residuals (see Tables 11 and 12) also indicate no serial correlation.
Table 11: Ljung–Box (LB) test p-values for residual autocorrelation of 3C 273 light curve.
Lag p-value
1 0.858400
2 0.443073
3 0.632991
4 0.677931
5 0.743433
6 0.778520
7 0.764781
8 0.755184
9 0.709967
10 0.579299
Table 12: Ljung–Box (LB) test p-values for residual autocorrelation of S4 0954+65 light curve.
Lag p-value
1 0.161701
2 0.293037
3 0.252978
4 0.361652
5 0.474434
6 0.565536
7 0.674115
8 0.630585
9 0.568174
10 0.649945
The resulting residual time series are therefore characteristic of a white noise sequence. Hence, we conclude that ARIMA(1,1,1)(1,1,1) and ARIMA(1,1,4)(1,1,4) have modelled the light curves of 3C 273 and S4 0954+65, respectively.
Figure 29: Autocorrelation (ACF) and Partial autocorrelation (PACF) plots for the light curve of 3C 273 (top) and the residuals (bottom) obtained after subtracting the best fit ARIMA(1,1,1)(1,1,1) model from the light curve data. The residual ACF/PACF plots show that the model fit has reduced most of the correlation present in the light curve data.
Figure 30: Autocorrelation (ACF) and Partial autocorrelation (PACF) plots for the differenced light curve of S4 0954+65 (top) and the residuals (bottom) obtained after subtracting the best fit ARIMA(1,1,4)(1,1,4) model from the light curve data. The residual ACF/PACF plots show that the model fit has reduced most of the correlation present in the light curve data.
@@ 5 Conclusions
In this paper, we presented a vectorized ARIMA–Nested Sampling framework and validated it on simulated time series data, recovering true model orders and parameters accurately. The framework was also applied to sunspot counts and to light curve data from Kepler and TESS. It successfully characterised the stochastic variability in these light curves and produced well-calibrated forecasts for the sunspot counts. We note that future improvements to the multi-step forecasts could be made by adopting a rolling window, which is a more accurate (although computationally expensive in our case) way of implementing ARIMA forecasting (Ndungi & Stanislavovich, 2025). Additionally, the forecasting can be made more robust within a Bayesian framework using a short time-step rolling forecast and consequently treating the obtained posterior distributions as priors for the next forecast iteration.
The sunspot analysis also revealed an important methodological finding. We find that statsmodels’ and pmdarima’s two independent Maximum Likelihood Estimation (MLE) implementations converged on ARIMA(3,0,3)(3,0,3) and ARIMA(3,0,2)(3,0,2) respectively via BIC minimization, while nested sampling selected ARIMA(9,0,1)(9,0,1) with maximum log-evidence. The superior forecasting performance of the nested sampling result on the held-out data demonstrates that Bayesian evidence can identify model structures that MLE-based selection systematically misses, owing to convergence failures in high-dimensional ARMA parameter spaces. This highlights a concrete practical advantage of the nested sampling approach beyond its theoretical appeal.
@@
The dominant computational bottleneck was the rejection sampling of valid ARMA coefficient vectors ($\phi_{a}$ and $\theta_{m}$) for higher orders ($p+q>8$) using the process discussed in Section 3. If very strong autocorrelation in the time-series is not expected a priori, then the “rejection sampling" process can be sped up by regularizing and tightening the priors using $\sigma^{\prime\prime}<1$ (Table 1). This is quantified in Appendix B (Table 14), which reports wall-clock runtimes for grid searches of increasing size and for different prior scales. A more powerful approach is offered by reparametrization of the ARMA(p,q)(p,q) coefficients to a set of $p+q$ partial autocorrelation coefficients (Barndorff-Nielsen & Schou, 1973). These coefficients are naturally constrained between 1-1 and $1$, and therefore a simple uniform prior, bounded between these two values, can be used. The vectorized implementation also offers a significant speed-up through parallelization of the nested sampling controlled by the $n_{\mathrm{delete}}$ factor. These computation times can be further improved through GPU-acceleration which our framework supports.
We note that, while the framework enables joint selection across all three ARIMA orders $p$, $d$ and $q$, the real-data applications in this paper fix $d$ a priori using classical stationarity tests (ADF and KPSS) and perform grid searches over $p$ and $q$ only. This was done for computational tractability and to ensure comparability of evidences on a common differenced series. Full three-dimensional grid searches including $d$ are demonstrated on simulated data (Section 4, Table LABEL:logp_d_table) and are entirely supported by the framework; we leave their application to real astronomical datasets with ambiguous stationarity as a direction for future work.
@@
Although limited to classic ARIMA models, the framework presented here can be extended to its hybrid variants such as Seasonal ARIMA (SARIMA), Seasonal ARIMA with exogenous variables (SARIMAX), Continuous ARIMA (CARIMA) and Autoregressive Fractionally Integrated Moving Average (ARFIMA) models (Feigelson et al., 2018). These extensions are efficient in modelling both periodic and stochastic variability exhibited in many astronomical time series, such as those of transiting exoplanets, eclipsing binaries or Active Galactic Nuclei (AGN).
@@
Finally, the investigation of potential physical interpretations behind the use of autoregressive modelling in time domain astronomy forms another compelling direction of future work. This question could be explored by applying our model selection procedure to different time-varying astrophysical systems. If systems of a particular class tend to favour similar model structures, then the ARIMA model may be representative of an underlying physical process rather than being a mere statistical description of the data.